Solve by substitution method and what is the solution set?
15x-10y=5
15y=-6+15x

Answer by faceoff57(108)

15x - 10y = 5
15y = 6 + 15x
15x - 10y = 5
15x = 10y + 5
x = 10/15y + 5/15
15y = 6 + 15(10/15y + 5/15)
15y = 6 + 10y + 5
15y = 10y + 11
15y - 10y = 11
5y = 11
y = 11/5 or 2 1/5
15x - 10(11/5) = 5
15x - 22 = 5
15x = 27
x = 27/15 or 1 12/15
(27/15, 11/5)